200MeV of energy obtained in the fission of one neucleus of 235^U. A reactor is generating 1000KW power. The rate of nuclear fission in the reactor is?
1
Expert's answer
2020-06-09T13:21:23-0400

The rate is the number of reactions per second. So we should divide the power by the energy in the fission of one nucleus. We should express the power in W and energy in J (1eV "\approx 1.60\cdot10^{-19}\,\mathrm{J}" )

"N = \dfrac{P}{E} = \dfrac{10^6\,\mathrm{W}}{2\cdot10^8\cdot (1.60\cdot10^{-19}\,\mathrm{J})} = 3.13\cdot10^{16}" fissions per second.